OpenStudy (anonymous):

The legs of a right triangles measure (m − 3) in. and (n + 3) in. Which is the length of square of its hypotenuse?

OpenStudy (anonymous):

please explain!

OpenStudy (anonymous):

\[(m-3)^{2}+(n-3)^{2}\] pythagoras theorem buddy

OpenStudy (anonymous):

\[(m-3)^2+(n+3)^2=h^2\]

OpenStudy (anonymous):

yes but how do you simplify it?

OpenStudy (anonymous):

answer choices : m2 + n2 − 6(m − n) + 18 m2 + n2 + 6(m − n) + 18 m2 + n2 + 6(m − n) + 9 m2 + n2 − 6(m

OpenStudy (anonymous):

square the each terms arrange the like terms together it's simple dude

OpenStudy (anonymous):

When you finish this could you please help me wiyh mine

OpenStudy (anonymous):

so m^2+n^2+(9+-9)? that doesnt get any of the answers

OpenStudy (anonymous):

dude \[(-3)^{2}=(3)^{2}=9\]

OpenStudy (anonymous):

oh -.-

OpenStudy (anonymous):

wait what??

OpenStudy (anonymous):

similarly \[(m+3)^{2}=m ^{2}+3^{2}+2*3*m\]

OpenStudy (anonymous):

\[(m-3)^2+(n+3)^2=h^2\] \[m^2-6m+9+n^2+6n+9\]